OpenStudy (anonymous):

The lengths of the corresponding sides of two similar trapezoids are in the ratio 3:5. If the area of the smaller trapezoid is 63sq inches, what is the area of the larger trapezoid?

OpenStudy (dumbcow):

side ratio is 3:5 area ratio is 3^2:5^2 = 9:25 A = 63*(25/9) A = 7*25 A = 175

OpenStudy (anonymous):

oh okay thanks!!!! <3